OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –6 3 3 6 Line 2 x y –3 1 3 3 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@Brostep0s

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

@MARC_ plz help

OpenStudy (anonymous):

Someone help me! plz!

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

i think it is C

OpenStudy (anonymous):

line 1=1/3 line 2=1/3

OpenStudy (anonymous):

plz don't leave like the others did

Directrix (directrix):

I'm not leaving. I am calculating the equations of the two lines over at http://www.mathportal.org/calculators/analytic-geometry/two-point-form-calculator.php

OpenStudy (anonymous):

i don't know if i should put general form or not

Directrix (directrix):

I agree with the slopes of the two lines being 1/3. Parallel lines have the same slope. The next question is if the two lines are the same.

OpenStudy (anonymous):

they are

OpenStudy (anonymous):

i think they are

Directrix (directrix):

The form does not really matter. I usually go with y = mx + b out of habit.

Directrix (directrix):

Check to see if the y-intercepts are the same.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it says y=1/3x+5

Directrix (directrix):

I got y-intercepts of 5 and 2. Did you get that?

Directrix (directrix):

Run that equation of a line calculator twice, once for each set of points.

OpenStudy (anonymous):

they are on the same line

Directrix (directrix):

i don't think so.

Directrix (directrix):

Click here and look at the graphs of the two lines. They have the same slope but not the same y-intercepts so they are different lines but do not intersect.

Directrix (directrix):

And, the answer is ? @TeenWolfGirl

OpenStudy (anonymous):

0 solution, A

Directrix (directrix):

Yes, that is what I got.

OpenStudy (anonymous):

thank you :)

Directrix (directrix):

Happy to help.